OpenStudy (anonymous):

what would be the solution of:

OpenStudy (anonymous):

\[15+or -i \sqrt{144}/6\]

OpenStudy (anonymous):

\[\frac{ 15+-isqrt{144} }{ 6}\]

OpenStudy (anonymous):

Solutions: -1 ± i -1 ± 2i -2 ± i -2 ± 2i

OpenStudy (anonymous):

I'm guessing the last one but idk

OpenStudy (btaylor):

Separate it into 2 fractions. You get \[\frac{15}{6}-\frac{i \sqrt{144}}{6}\]This simplifies into 2.5-12i/6 = 2.5-2i

OpenStudy (btaylor):

Are you sure it is 15?

OpenStudy (anonymous):

Oh no im wrong! It's suppose to be a 6

OpenStudy (anonymous):

wow thanks for pointing that out

OpenStudy (btaylor):

Oh. Then it would be 6/6 = 1 +- 2i, so I'd pick the second answer. Are you sure it isn't -6, since all you answers are -1 or -2?

OpenStudy (anonymous):

It was like this \[\frac{ -(-6) squrt (-6)^2-4(3)(15) }{ 6}\]

OpenStudy (anonymous):

Then I changed it to that @BTaylor

OpenStudy (btaylor):

Alright, it looks like you did it correctly. So I assume the original polynomial was: 3x^2-6x+15?

OpenStudy (anonymous):

Yep thanks for the help :)

OpenStudy (btaylor):

YW.